OpenStudy (anonymous):

Given that f(x)=sqrt{2 + x} and g(x)=sqrt{2 - x}, find formulas for the following functions, and their domains. In each case, enter the domain using interval notation. (a) f+g= and its domain is

OpenStudy (anonymous):

Have you searched the web? I havn't studied this yet /:

OpenStudy (anonymous):

I have.. I can't find anything similar :/

OpenStudy (anonymous):

@qpHalcy0n can you help me?

OpenStudy (anonymous):

@jim_thompson5910 can you please help? :)

jimthompson5910 (jim_thompson5910):

f(x) + g(x) literally means "add the functions f(x) and g(x)"

jimthompson5910 (jim_thompson5910):

but since f(x) = sqrt(2+x) and g(x) = sqrt(2-x), we know that f(x) + g(x) = sqrt(2+x) + sqrt(2-x)

OpenStudy (anonymous):

that's what I got.. But, I wasn't sure if f(x) + g(x) = sqrt(2+x) + sqrt(2-x) was correct or not.. how would I find the domain?

jimthompson5910 (jim_thompson5910):

yeah it seems too simple that it looks like a trick question, but it's not

jimthompson5910 (jim_thompson5910):

the domain is the set of allowable inputs, or x values

jimthompson5910 (jim_thompson5910):

remember that you can't take the square root of a negative number, so 2+x >= 0 and 2-x >= 0

jimthompson5910 (jim_thompson5910):

those two inequalities become x >= -2 and x <= 2 which combine to -2 <= x <= 2

OpenStudy (anonymous):

so, in interval notation.. it would be (-inf,-2)U(2,inf) ?

jimthompson5910 (jim_thompson5910):

no, [-2, 2]

jimthompson5910 (jim_thompson5910):

that represents \(\Large -2 \le x \le 2\)

OpenStudy (anonymous):

oh okay... and would this be correct? f(x) - g(x) = sqrt(2+x) - sqrt(2-x)

jimthompson5910 (jim_thompson5910):

yes that's 100% correct

jimthompson5910 (jim_thompson5910):

the domain is the same because the radicands (the stuff in the square roots) are the same

OpenStudy (anonymous):

f(x) * g(x) = sqrt(2+x) * sqrt(2-x) ?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

domains are the same (for the same reason explained above)

jimthompson5910 (jim_thompson5910):

domain is*

OpenStudy (anonymous):

f(x)/g(x) = sqrt(2+x)/sqrt(2-x) and same domain?

jimthompson5910 (jim_thompson5910):

the first part is correct, but the domain will be slightly different

jimthompson5910 (jim_thompson5910):

now we have to avoid dividing by zero, so 2-x can't be zero 2-x = 0 x = 2 so if x = 2, then 2-x is zero. So we toss out x = 2 from the domain So the domain is now [-2, 2)

OpenStudy (anonymous):

so now what is the difference between f+g and fog(x)?

jimthompson5910 (jim_thompson5910):

f+g means "add f and g" fog(x) means "start with f(x), and plug in g(x) as the input"

jimthompson5910 (jim_thompson5910):

fog(x) is the same as f(g(x))

jimthompson5910 (jim_thompson5910):

\[\Large f(x) = \sqrt{2+x}\] \[\Large f(g(x)) = \sqrt{2+g(x)}\] \[\Large f(g(x)) = \sqrt{2+\sqrt{2-x}}\]

OpenStudy (anonymous):

well it's with different numbers, I just wanted to know the difference.

jimthompson5910 (jim_thompson5910):

i gotcha, do you see how I'm getting that?

OpenStudy (anonymous):

f(g(x))=2+sqrt(2-x)?

jimthompson5910 (jim_thompson5910):

f(x) = sqrt(2+x) and g(x) = sqrt(2-x) correct?

OpenStudy (anonymous):

yes.

jimthompson5910 (jim_thompson5910):

Then \[\Large f(g(x)) = \sqrt{2+\sqrt{2-x}}\] or \[\Large f\circ g(x) = \sqrt{2+\sqrt{2-x}}\]

OpenStudy (anonymous):

can you draw that? cause its not showing up correct.,. and im a lil confused

jimthompson5910 (jim_thompson5910):